In eight throws of a dice ,5 or 6 is considered success.Find mean number of successes and standard deviation.
1
Expert's answer
2011-06-07T05:08:30-0400
It is binomial distibution with n=8 and p=P(success)= P(5 or 6)=1/6 + 1/6 =1/3
So q=1-p=2/3. Then the mean number of success is n*p=8*1/3=8/3=2.667
and standard deviation is
s=sqrt(n*p*q)=sqrt(8 * 1/3 * 2/3)=4/3=1.333.